# Bootstrapping the critical behavior of multi-matrix models

### v1

@@ B.2 The moments
The following relations were found by solving the SDE in Mathematica:
@@
$m_{AABB}$ $=\frac{1-m_{2}}{4g}$
$m_{ABAB}$ $=\frac{4gm_{2}^{2}+m_{2}-1}{4g}$
@@
$m_{AAAABB}$ $=\frac{-4g^{2}m_{2}^{3}-4gm_{2}^{2}+8gm_{2}+m_{2}-1}{16g^{2}}$
$m_{AAABAB}$ $=\frac{-4g^{2}m_{2}^{3}+4gm_{2}^{2}+m_{2}-1}{16g^{2}}$
$m_{AABAAB}$ $=\frac{12g^{2}m_{2}^{3}+4gm_{2}^{2}+m_{2}-1}{16g^{2}}$
@@
$m_{AAAAAABB}$ $=\frac{-16g^{3}m_{2}^{4}+36g^{2}m_{2}^{2}+12gm_{2}^{2}-24gm_{2}+8g-m_{2}+1}{64g^{3}}$
$m_{AAAAABAB}$ $=\frac{16g^{3}m_{2}^{4}+32g^{2}m_{2}^{3}-28g^{2}m_{2}^{2}-8gm_{2}+4g-m_{2}+1}{64g^{3}}$
$m_{AAAABAAB}$ $=\frac{16g^{3}m_{2}^{4}-16g^{2}m_{2}^{3}+20g^{2}m_{2}^{2}-4gm_{2}^{2}-m_{2}+1}{64g^{3}}$
$m_{AAAABBBB}$ $=\frac{-16g^{3}m_{2}^{4}+36g^{2}m_{2}^{2}+8gm_{2}^{2}-16gm_{2}+4g-m_{2}+1}{64g^{3}}$
$m_{AAABAAAB}$ $=\frac{-48g^{3}m_{2}^{4}-16g^{2}m_{2}^{3}+20g^{2}m_{2}^{2}-8gm_{2}^{2}+8gm_{2}-4g-m_{2}+1}{64g^{3}}$
$m_{AAABABBB}$ $=\frac{16g^{3}m_{2}^{4}+16g^{2}m_{2}^{3}-12g^{2}m_{2}^{2}-4gm_{2}^{2}-m_{2}+1}{64g^{3}}$
$m_{AAABBABB}$ $=\frac{16g^{3}m_{2}^{4}+4g^{2}m_{2}^{2}-8gm_{2}+4g-m_{2}+1}{64g^{3}}$
$m_{AABABABB}$ $=\frac{-16g^{3}m_{2}^{4}-28g^{2}m_{2}^{2}-4gm_{2}^{2}-8gm_{2}+8g-m_{2}+1}{64g^{3}}$
$m_{AABABBAB}$ $=\frac{-16g^{3}m_{2}^{4}-32g^{2}m_{2}^{3}+4g^{2}m_{2}^{2}-8gm_{2}^{2}+4g-m_{2}+1}{64g^{3}}$
$m_{AABBAABB}$ $=\frac{16g^{3}m_{2}^{4}+16g^{2}m_{2}^{3}-12g^{2}m_{2}^{2}+8gm_{2}^{2}-24gm_{2}+12g-m_{2}+1}{64g^{3}}$
$m_{ABABABAB}$ $=\frac{16g^{3}m_{2}^{4}-16g^{2}m_{2}^{3}-44g^{2}m_{2}^{2}-8gm_{2}^{2}-8gm_{2}+12g-m_{2}+1}{64g^{3}}$
@@ C.2 The moments
$m_{AABB}$ $=\frac{-1+m_{2}+gm_{2}+gm_{4}}{3g}$
$m_{ABAB}$ $=-\frac{-1+m_{2}-2gm_{2}+gm_{4}}{3g)}$
@@
$m_{AAAABB}$ $=\frac{-2-2g+2m_{2}-2gm_{2}-g^{2}m_{2}+3g^{2}m_{2}^{2}+2gm_{4}+8g^{2}m_{4}}{12g^{2}}$
$m_{AAABAB}$ $=\frac{2-2m_{2}+4gm_{2}+3g^{2}m_{2}+3g^{2}m_{2}^{2}-2gm_{4}}{12g^{2}}$
$m_{AABAAB}$ $=\frac{-2+2m_{2}-4gm_{2}+3g^{2}m_{2}+3g^{2}m_{2}^{2}+2gm_{4}}{12g^{2}}$
@@ D.2 The moments
$m_{AABB}$ $=\frac{1-m_{2}+gm_{2}+gm_{4}}{3g}$
$m_{ABAB}$ $=\frac{1-m_{2}-2gm_{2}+gm_{4}}{3g}$
@@
$m_{AAAABB}$ $=\frac{-2+2g+2m_{2}+2gm_{2}-g^{2}m_{2}+3g^{2}m_{2}^{2}-2gm_{4}+8g^{2}m_{4}}{12g^{2}}$
$m_{AAABAB}$ $=\frac{-2+2m_{2}+4gm_{2}-3g^{2}m_{2}-3g^{2}m_{2}^{2}-2gm_{4}}{12g^{2}}$
$m_{AABAAB}$ $=\frac{-2+2m_{2}+4gm_{2}+3g^{2}m_{2}+3g^{2}m_{2}^{2}-2gm_{4}}{12g^{2}}$
@@ E.1 The SDE
$A:0$ $=1-m_{2}-gm_{3}-2gm_{1,1,1}$
$B:0$ $=-m_{1,1}-3gm_{2,1}$
$C:0$ $=-m_{1,1}-3gm_{2,1}$
$A^{2}:0$ $=2m_{1}-2gm_{2,1,1}-m_{3}-gm_{4}$
$AB:0$ $=m_{1}-gm_{2,1,1}-gm_{1,1,0,1,0,1}-m_{2,1}-gm_{3,1}$
$AC:0$ $=m_{1}-gm_{2,1,1}-gm_{1,1,0,1,0,1}-m_{2,1}-gm_{3,1}$
$BA:0$ $=m_{1}-gm_{2,1,1}-gm_{1,1,0,1,0,1}-m_{2,1}-gm_{3,1}$
$B^{2}:0$ $=-m_{2,1}-2gm_{3,1}-gm_{2,2}$
$BC:0$ $=-gm_{2,1,1}-gm_{1,1,0,1,1}-m_{1,1,1}-gm_{2,2}$
$CA:0$ $=m_{1}-gm_{2,1,1}-gm_{1,1,0,1,0,1}-m_{2,1}-gm_{3,1}$
$CB:0$ $=-gm_{2,1,1}-gm_{1,1,0,1,1}-m_{1,1,1}-gm_{2,2}$
$C^{2}:0$ $=-m_{2,1}-2gm_{3,1}-gm_{2,2}$
$A^{3}:0$ $=m_{1}^{2}-gm_{5}-2gm_{3,1,1}+2m_{2}-m_{4}$
$A^{2}B:0$ $=m_{1}^{2}-gm_{4,1}-gm_{2,2,1}-gm_{2,1,1,0,1}+m_{1,1}-m_{3,1}$
$A^{2}C:0$ $=m_{1}^{2}-gm_{4,1}-gm_{2,2,1}-gm_{2,1,1,0,1}+m_{1,1}-m_{3,1}$
$ABA:0$ $=-gm_{4,1}-2gm_{1,1,0,1,1,1}+2m_{1,1}-m_{3,1}$
$AB^{2}:0$ $=-gm_{3,2}-gm_{3,1,1}-gm_{2,1,0,1,0,1}+m_{2}-m_{2,2}$
$ABC:0$ $=-m_{2,1,1}-gm_{3,1,1}-gm_{2,1,1,0,1}-gm_{1,1,0,1,1,1}+m_{1,1}$
$ACA:0$ $=-gm_{4,1}-2gm_{1,1,0,1,1,1}+2m_{1,1}-m_{3,1}$
$ACB:0$ $=-m_{2,1,1}-gm_{3,1,1}-gm_{2,1,1,0,1}-gm_{1,1,0,1,1,1}+m_{1,1}$
$AC^{2}:0$ $=-gm_{3,2}-gm_{3,1,1}-gm_{2,1,0,1,0,1}+m_{2}-m_{2,2}$
$BA^{2}:0$ $=m_{1}^{2}-gm_{4,1}-gm_{2,2,1}-gm_{2,1,1,0,1}+m_{1,1}-m_{3,1}$
@@ E.2 The moments
$m_{2}$ $=\frac{-m_{1}-2gm_{2}}{g}$
$m_{3}$ $=\frac{g+m_{1}-6gm_{1}^{2}+2gm_{2}-18g^{2}m_{1}m_{2}}{3g^{2}}$
$m_{AAB}$ $=-\frac{m_{2}}{3g}$
$m_{ABC}$ $=\frac{g+m_{1}+3gm_{1}^{2}+2gm_{2}+9g^{2}m_{1}m_{2}}{3g^{2}}$
$m_{4}$ $=\frac{-7g-7m_{1}+24g^{2}m_{1}+48gm_{1}^{2}-36g^{2}m_{1}^{3}-18gm_{2}-18g^{3}m_{2}+144g^{2}m_{1}m_{2}-108g^{3}m_{1}^{2}m_{2}}{27g^{3}}$
$m_{AAAB}$ $=\frac{2g+2m_{1}+6g^{2}m_{1}+3gm_{1}^{2}-9g^{2}m_{1}^{3}+9gm_{2}+9g^{3}m_{2}+9g^{2}m_{1}m_{2}-27g^{3}m_{1}^{2}m_{2}}{27g^{3}}$
$m_{AABB}$ $=\frac{-4g-4m_{1}-12g^{2}m_{1}-6gm_{1}^{2}+18g^{2}m_{1}^{3}-9gm_{2}-18g^{3}m_{2}-18g^{2}m_{1}m_{2}+54g^{3}m_{1}^{2}m_{2}}{27g^{3}}$
$m_{AABC}$ $=-\frac{g+m_{1}-15g^{2}m_{1}-3gm_{1}^{2}-18g^{2}m_{1}^{3}-9g^{3}m_{2}-9g^{2}m_{1}m_{2}-54g^{3}m_{1}^{2}m_{2}}{27g^{3}}$
$m_{ABAB}$ $=\frac{-4g-4m_{1}-3g^{2}m_{1}-24gm_{1}^{2}-36g^{2}m_{1}^{3}-9gm_{2}+9g^{3}m_{2}-72g^{2}m_{1}m_{2}-108g^{3}m_{1}^{2}m_{2}}{27g^{3}}$
$m_{ABAC}$ $=\frac{-g-m_{1}+6g^{2}m_{1}-6gm_{1}^{2}-9g^{2}m_{1}^{3}-18g^{3}m_{2}-18g^{2}m_{1}m_{2}-27g^{3}m_{1}^{2}m_{2}}{27g^{3}}$
$m_{5}$ $=\frac{1}{81g^{4}}(11g+11m_{1}-192g^{2}m_{1}-150gm_{1}^{2}+9g^{3}m_{1}^{2}+180g^{2}m_{1}^{3}+30gm_{2}-180g^{3}m_{2}$
$-450g^{2}m_{1}m_{2}-270g^{4}m_{1}m_{2}+540g^{3}m_{1}^{2}m_{2})$
$m_{AAAAB}$ $=\frac{1}{81g^{4}}(-4g-4m_{1}+6g^{2}m_{1}+3gm_{1}^{2}+36g^{3}m_{1}^{2}+45g^{2}m_{1}^{3}-15gm_{2}+9g^{3}m_{2}$
$+9g^{2}m_{1}m_{2}+54g^{4}m_{1}m_{2}+135g^{3}m_{1}^{2}m_{2})$
$m_{AAABB}$ $=\frac{1}{81g^{4}}(2g+2m_{1}-30g^{2}m_{1}-15gm_{1}^{2}-72g^{3}m_{1}^{2}-63g^{2}m_{1}^{3}+3gm_{2}-45g^{3}m_{2}$
$-45g^{2}m_{1}m_{2}-108g^{4}m_{1}m_{2}-189g^{3}m_{1}^{2}m_{2})$
$m_{AAABC}$ $=\frac{1}{81g^{4}}(5g+5m_{1}-21g^{2}m_{1}+3gm_{1}^{2}+36g^{3}m_{1}^{2}-36g^{2}m_{1}^{3}+12gm_{2}-45g^{3}m_{2}+9g^{2}m_{1}m_{2}$
$+135g^{4}m_{1}m_{2}-108g^{3}m_{1}^{2}m_{2})$
$m_{AABAB}$ $=\frac{1}{81g^{4}}(2g+2m_{1}-3g^{2}m_{1}+12gm_{1}^{2}+9g^{3}m_{1}^{2}+18g^{2}m_{1}^{3}+3gm_{2}+9g^{3}m_{2}+36g^{2}m_{1}m_{2}$
$+54g^{4}m_{1}m_{2}+54g^{3}m_{1}^{2}m_{2})$
$m_{AABAC}$ $=\frac{1}{81g^{4}}(5g+5m_{1}+6g^{2}m_{1}+30gm_{1}^{2}+36g^{3}m_{1}^{2}+45g^{2}m_{1}^{3}+12gm_{2}-18g^{3}m_{2}+90g^{2}m_{1}m_{2}$
$-27g^{4}m_{1}m_{2}+135g^{3}m_{1}^{2}m_{2})$
$m_{AABBC}$ $=-\frac{1}{81g^{4}}(g+m_{1}+12g^{2}m_{1}+6gm_{1}^{2}-63g^{3}m_{1}^{2}+9g^{2}m_{1}^{3}+6gm_{2}-9g^{3}m_{2}+18g^{2}m_{1}m_{2}$
$-54g^{4}m_{1}m_{2}+27g^{3}m_{1}^{2}m_{2})$
$m_{AABCB}$ $=\frac{1}{81g^{4}}(-g-m_{1}-12g^{2}m_{1}-6gm_{1}^{2}-18g^{3}m_{1}^{2}-9g^{2}m_{1}^{3}-6gm_{2}+36g^{3}m_{2}-18g^{2}m_{1}m_{2}$
$-108g^{4}m_{1}m_{2}-27g^{3}m_{1}^{2}m_{2})$
$m_{ABABC}$ $=\frac{1}{81g^{4}}(-g-m_{1}-12g^{2}m_{1}-6gm_{1}^{2}-18g^{3}m_{1}^{2}-9g^{2}m_{1}^{3}-6gm_{2}+63g^{3}m_{2}-18g^{2}m_{1}m_{2}$
$-27g^{4}m_{1}m_{2}-27g^{3}m_{1}^{2}m_{2})$

### v2

@@ B.2 The moments
The following relations were found by solving some large set of SDE in Mathematica:
@@
$m_{2,2}$ $=\frac{1-m_{2}}{4g}$
$m_{1,1,1,1}$ $=\frac{4gm_{2}^{2}+m_{2}-1}{4g}$
@@
$m_{4,2}$ $=\frac{-4g^{2}m_{2}^{3}-4gm_{2}^{2}+8gm_{2}+m_{2}-1}{16g^{2}}$
$m_{3,1,1,1}$ $=\frac{-4g^{2}m_{2}^{3}+4gm_{2}^{2}+m_{2}-1}{16g^{2}}$
$m_{2,1,2,1}$ $=\frac{12g^{2}m_{2}^{3}+4gm_{2}^{2}+m_{2}-1}{16g^{2}}$
@@
$m_{6,2}$ $=\frac{-16g^{3}m_{2}^{4}+36g^{2}m_{2}^{2}+12gm_{2}^{2}-24gm_{2}+8g-m_{2}+1}{64g^{3}}$
$m_{5,1,1,1}$ $=\frac{16g^{3}m_{2}^{4}+32g^{2}m_{2}^{3}-28g^{2}m_{2}^{2}-8gm_{2}+4g-m_{2}+1}{64g^{3}}$
$m_{4,1,2,1}$ $=\frac{16g^{3}m_{2}^{4}-16g^{2}m_{2}^{3}+20g^{2}m_{2}^{2}-4gm_{2}^{2}-m_{2}+1}{64g^{3}}$
$m_{4,4}$ $=\frac{-16g^{3}m_{2}^{4}+36g^{2}m_{2}^{2}+8gm_{2}^{2}-16gm_{2}+4g-m_{2}+1}{64g^{3}}$
$m_{3,1,3,1}$ $=\frac{-48g^{3}m_{2}^{4}-16g^{2}m_{2}^{3}+20g^{2}m_{2}^{2}-8gm_{2}^{2}+8gm_{2}-4g-m_{2}+1}{64g^{3}}$
$m_{3,1,1,3}$ $=\frac{16g^{3}m_{2}^{4}+16g^{2}m_{2}^{3}-12g^{2}m_{2}^{2}-4gm_{2}^{2}-m_{2}+1}{64g^{3}}$
$m_{3,2,1,2}$ $=\frac{16g^{3}m_{2}^{4}+4g^{2}m_{2}^{2}-8gm_{2}+4g-m_{2}+1}{64g^{3}}$
$m_{2,1,1,1,1,2}$ $=\frac{-16g^{3}m_{2}^{4}-28g^{2}m_{2}^{2}-4gm_{2}^{2}-8gm_{2}+8g-m_{2}+1}{64g^{3}}$
$m_{2,1,1,2,1,1}$ $=\frac{-16g^{3}m_{2}^{4}-32g^{2}m_{2}^{3}+4g^{2}m_{2}^{2}-8gm_{2}^{2}+4g-m_{2}+1}{64g^{3}}$
$m_{2,2,2,2}$ $=\frac{16g^{3}m_{2}^{4}+16g^{2}m_{2}^{3}-12g^{2}m_{2}^{2}+8gm_{2}^{2}-24gm_{2}+12g-m_{2}+1}{64g^{3}}$
$m_{1,1,1,1,1,1,1,1}$ $=\frac{16g^{3}m_{2}^{4}-16g^{2}m_{2}^{3}-44g^{2}m_{2}^{2}-8gm_{2}^{2}-8gm_{2}+12g-m_{2}+1}{64g^{3}}$
@@ C.2 The moments
The following relations were found by solving some large set of SDE in Mathematica:
$m_{2,2}$ $=\frac{-1+m_{2}+gm_{2}+gm_{4}}{3g}$
$m_{1,1,1,1}$ $=-\frac{-1+m_{2}-2gm_{2}+gm_{4}}{3g)}$
@@
$m_{4,2}$ $=\frac{-2-2g+2m_{2}-2gm_{2}-g^{2}m_{2}+3g^{2}m_{2}^{2}+2gm_{4}+8g^{2}m_{4}}{12g^{2}}$
$m_{3,1,1,1}$ $=\frac{2-2m_{2}+4gm_{2}+3g^{2}m_{2}+3g^{2}m_{2}^{2}-2gm_{4}}{12g^{2}}$
$m_{2,1,2,1}$ $=\frac{-2+2m_{2}-4gm_{2}+3g^{2}m_{2}+3g^{2}m_{2}^{2}+2gm_{4}}{12g^{2}}$
@@ D.2 The moments
The following relations were found by solving some large set of SDE in Mathematica:
$m_{2,2}$ $=\frac{1-m_{2}+gm_{2}+gm_{4}}{3g}$
$m_{1,1,1,1}$ $=\frac{1-m_{2}-2gm_{2}+gm_{4}}{3g}$
@@
$m_{4,2}$ $=\frac{-2+2g+2m_{2}+2gm_{2}-g^{2}m_{2}+3g^{2}m_{2}^{2}-2gm_{4}+8g^{2}m_{4}}{12g^{2}}$
$m_{3,1,1,1}$ $=\frac{-2+2m_{2}+4gm_{2}-3g^{2}m_{2}-3g^{2}m_{2}^{2}-2gm_{4}}{12g^{2}}$
$m_{2,1,2,1}$ $=\frac{-2+2m_{2}+4gm_{2}+3g^{2}m_{2}+3g^{2}m_{2}^{2}-2gm_{4}}{12g^{2}}$
@@ E.1 The SDE
The following are some examples of the SDE for this model:
$A:0$ $=1-m_{AA}-gm_{AAA}-2gm_{ABC}$
$B:0$ $=-m_{AB}-3gm_{AAB}$
$C:0$ $=-m_{AB}-3gm_{AAB}$
$A^{2}:0$ $=2m_{A}-2gm_{AABC}-m_{AAA}-gm_{AAAA}$
$AB:0$ $=m_{A}-gm_{AABC}-gm_{ABAC}-m_{AAB}-gm_{AAAB}$
$AC:0$ $=m_{A}-gm_{AABC}-gm_{ABAC}-m_{AAB}-gm_{AAAB}$
$BA:0$ $=m_{A}-gm_{AABC}-gm_{ABAC}-m_{AAB}-gm_{AAAB}$
$B^{2}:0$ $=-m_{AAB}-2gm_{AAAB}-gm_{AABB}$
$BC:0$ $=-gm_{AABC}-gm_{ABAB}-m_{ABC}-gm_{AABB}$
$CA:0$ $=m_{A}-gm_{AABC}-gm_{ABAC}-m_{AAB}-gm_{AAAB}$
$CB:0$ $=-gm_{AABC}-gm_{ABAB}-m_{ABC}-gm_{AABB}$
$C^{2}:0$ $=-m_{AAB}-2gm_{AAAB}-gm_{AABB}$
$A^{3}:0$ $=m_{A}^{2}-gm_{AAAAA}-2gm_{AAABC}+2m_{AA}-m_{AAAA}$
$A^{2}B:0$ $=m_{A}^{2}-gm_{AAAAB}-gm_{AABBC}-gm_{AABCB}+m_{AB}-m_{AAAB}$
$A^{2}C:0$ $=m_{A}^{2}-gm_{AAAAB}-gm_{AABBC}-gm_{AABCB}+m_{AB}-m_{AAAB}$
$ABA:0$ $=-gm_{AAAAB}-2gm_{ABABC}+2m_{AB}-m_{AAAB}$
$AB^{2}:0$ $=-gm_{AAABB}-gm_{AAABC}-gm_{AABAC}+m_{AA}-m_{AABB}$
$ABC:0$ $=-m_{AABC}-gm_{AAABC}-gm_{AABCB}-gm_{ABABC}+m_{AB}$
$ACA:0$ $=-gm_{AAAAB}-2gm_{ABABC}+2m_{AB}-m_{AAAB}$
$ACB:0$ $=-m_{AABC}-gm_{AAABC}-gm_{AABCB}-gm_{ABABC}+m_{AB}$
$AC^{2}:0$ $=-gm_{AAABB}-gm_{AAABC}-gm_{AABAC}+m_{AA}-m_{AABB}$
$BA^{2}:0$ $=m_{A}^{2}-gm_{AAAAB}-gm_{AABBC}-gm_{AABCB}+m_{AB}-m_{AAAB}$
@@ E.2 The moments
The following relations were found by solving some large set of SDE in Mathematica:
$m_{AB}=$ $\frac{-m_{1}-gm_{2}}{2g}$
$m_{AAA}=$ $\frac{1+3m_{1}^{2}-m_{2}+9gm_{1}m_{2}}{3g}$
$m_{AAB}=$ $\frac{m_{1}+gm_{2}}{6g^{2}}$
$m_{ABC}=$ $\frac{2-3m_{1}^{2}-2m_{2}-9gm_{1}m_{2}}{6g}$
$m_{AAAA}=$ $\frac{-7g+2m_{1}+33g^{2}m_{1}-24gm_{1}^{2}+18g^{2}m_{1}^{3}+9gm_{2}+9g^{3}m_{2}-72g^{2}m_{1}m_{2}+54g^{3}m_{1}^{2}m_{2}}{27g^{3}}$
$m_{AAAB}=$ $\frac{4g-5m_{1}+3g^{2}m_{1}-3gm_{1}^{2}+9g^{2}m_{1}^{3}-9gm_{2}-9g^{3}m_{2}-9g^{2}m_{1}m_{2}+27g^{3}m_{1}^{2}m_{2}}{54g^{3}}$
$m_{AABB}=$ $\frac{-8g+m_{1}-6g^{2}m_{1}+6gm_{1}^{2}-18g^{2}m_{1}^{3}+9gm_{2}+18g^{3}m_{2}+18g^{2}m_{1}m_{2}-54g^{3}m_{1}^{2}m_{2}}{54g^{3}}$
$m_{AABC}=$ $\frac{-2g-2m_{1}+21g^{2}m_{1}-3gm_{1}^{2}-18g^{2}m_{1}^{3}-9g^{3}m_{2}-9g^{2}m_{1}m_{2}-54g^{3}m_{1}^{2}m_{2}}{54g^{3}}$
$m_{ABAB}=$ $\frac{-8g+m_{1}-15g^{2}m_{1}+24gm_{1}^{2}+36g^{2}m_{1}^{3}+9gm_{2}-9g^{3}m_{2}+72g^{2}m_{1}m_{2}+108g^{3}m_{1}^{2}m_{2}}{54g^{3}}$
$m_{ABAC}=$ $\frac{-2g-2m_{1}+30g^{2}m_{1}+6gm_{1}^{2}+9g^{2}m_{1}^{3}+18g^{3}m_{2}+18g^{2}m_{1}m_{2}+27g^{3}m_{1}^{2}m_{2}}{54g^{3}}$
$m_{AAAAA}=$ $\frac{1}{81g^{4}}\left(11g-4m_{1}-102g^{2}m_{1}+75gm_{1}^{2}+144g^{3}m_{1}^{2}-90g^{2}m_{1}^{3}-15gm_{2}+90g^{3}m_{2}\right.$
$\left.+225g^{2}m_{1}m_{2}+135g^{4}m_{1}m_{2}-270g^{3}m_{1}^{2}m_{2}\right)$
$m_{AAAAB}=$ $\frac{1}{162g^{4}}\left(-8g+7m_{1}+3g^{2}m_{1}-3gm_{1}^{2}+18g^{3}m_{1}^{2}-45g^{2}m_{1}^{3}+15gm_{2}\right.$
$\left.-9g^{3}m_{2}-9g^{2}m_{1}m_{2}-54g^{4}m_{1}m_{2}-135g^{3}m_{1}^{2}m_{2}\right)$
$m_{AAABB}=$ $\frac{1}{162g^{4}}\left(4g+m_{1}-15g^{2}m_{1}+15gm_{1}^{2}-36g^{3}m_{1}^{2}+63g^{2}m_{1}^{3}-3gm_{2}\right.$
$\left.+45g^{3}m_{2}+45g^{2}m_{1}m_{2}+108g^{4}m_{1}m_{2}+189g^{3}m_{1}^{2}m_{2}\right)$
$m_{AAABC}=$ $\frac{1}{162g^{4}}\left(10g-2m_{1}+3g^{2}m_{1}-3gm_{1}^{2}-63g^{3}m_{1}^{2}+36g^{2}m_{1}^{3}-12gm_{2}\right.$
$\left.+45g^{3}m_{2}-9g^{2}m_{1}m_{2}-135g^{4}m_{1}m_{2}+108g^{3}m_{1}^{2}m_{2}\right)$
$m_{AABAB}=$ $\frac{1}{162g^{4}}\left(4g+m_{1}-15g^{2}m_{1}-12gm_{1}^{2}-36g^{3}m_{1}^{2}-18g^{2}m_{1}^{3}-3gm_{2}\right.$
$\left.-9g^{3}m_{2}-36g^{2}m_{1}m_{2}-54g^{4}m_{1}m_{2}-54g^{3}m_{1}^{2}m_{2}\right)$
$m_{AABAC}=$ $\frac{1}{162g^{4}}\left(10g-2m_{1}+30g^{2}m_{1}-30gm_{1}^{2}+99g^{3}m_{1}^{2}-45g^{2}m_{1}^{3}-12gm_{2}\right.$
$\left.+18g^{3}m_{2}-90g^{2}m_{1}m_{2}+27g^{4}m_{1}m_{2}-135g^{3}m_{1}^{2}m_{2}\right)$